OpenStudy (anonymous):

How do you write the quadratic equation which has the roots: 2+√ 3 and 2-√ 3

jimthompson5910 (jim_thompson5910):

\[\Large x = 2 + \sqrt{3} \ \text{or} \ x = 2 - \sqrt{3}\] \[\Large x = 2 \pm \sqrt{3}\] \[\Large x - 2 = \pm \sqrt{3}\] See where to go from here?

OpenStudy (anonymous):

no can you show me please

jimthompson5910 (jim_thompson5910):

\[\Large x = 2 + \sqrt{3} \ \text{or} \ x = 2 - \sqrt{3}\] \[\Large x = 2 \pm \sqrt{3}\] \[\Large x - 2 = \pm \sqrt{3}\] \[\Large (x - 2)^2 = 3\] \[\Large (x - 2)^2 - 3 = 0\] From here, you expand and simplify.

OpenStudy (anonymous):

oh okay thank you I understand now!

jimthompson5910 (jim_thompson5910):

yw